danielaaxrosas89:

How do I find the values of x and y? Plus, how do I write my answer in simplest form. IMG_5801.jpg

danielaaxrosas89:

1 attachment
iGreen:

Are you familiar with sine, cosine, and tangent?

danielaaxrosas89:

Yes... but I don’t know how to use those formulas

iGreen:

Do you remember what they equal?

danielaaxrosas89:

90 degrees?

iGreen:

No... \(\sf sine=\dfrac{opposite}{hypotenuse}\) \(\sf cosine=\dfrac{adjacent}{hypotenuse}\) \(\sf tangent=\dfrac{opposite}{adjacent}\)

iGreen:

@danielaaxrosas89

danielaaxrosas89:

oh my bad lol

iGreen:

So which one can you use to calculate the sides?

iGreen:

@danielaaxrosas89

danielaaxrosas89:

probably tangent?

iGreen:

No, tangent is opposite over adjacent, that's the only one we can't use because we have two unknowns(x and y).

iGreen:

We know the hypotenuse and the angle.

iGreen:

@danielaaxrosas89

danielaaxrosas89:

So cosine?

iGreen:

Sorry, I had to go out for a few hours...